# The Cosmic Dance: Observational Detection of Coherent Spin in Galaxy Clusters

### v1

@@ Abstract
The spin of galaxy clusters encodes key information about their formation, dynamics, and the influence of large-scale structure. However, whether clusters possess statistically significant spin and how to measure it observationally remain open questions. Here, we present the first observational statistical detection of coherent spin in galaxy clusters, by using a sample of 2,170 systems with $M>10^{14}\,M_{\odot}$ selected from a publicly available group catalog based on the SDSS galaxy. Cluster spin is quantified by identifying the orientation in the projected plane that maximizes the redshift difference ($\Delta Z_{\rm max}$) between member galaxies in two regions divided by a trial axis. We find strong statistical evidence for coherent rotation, with the observed $\Delta Z_{\rm max}$ distribution significantly exceeding randomized controls (nearly $\sim 150\sigma$ confidence at $\sim 380{\,\rm km\,s^{-1}}$), especially in richer clusters ($N_{\mathrm{gal}}>10$, up to $\sim 300\sigma$). Stacked visualizations confirm the spatial segregation of redshifted and blueshifted galaxies across the rotation axis. The radial profile of the rotational velocity indicates that it increases as a function of radius. The cluster rotation speed increases with mass, from $\sim 330~\mathrm{km\,s}^{-1}$ at $10^{14}M_{\odot}$ to $\sim 800~\mathrm{km\,s}^{-1}$ at $10^{15}M_{\odot}$. Additionally, cluster spin tends to align parallel with the central galaxy spin and perpendicular to the nearest cosmic filament, particularly in richer systems. These results reveal significant coherent spin in galaxy clusters, shaped by both internal dynamics and large-scale structure.
@@ 1 Introduction
Simulations have also provided valuable insights into cluster spin. Baldi et al. (2019) used MUSIC simulations and kinetic Sunyaev-Zel’dovich (kSZ) maps to detect rotational signals averaging 23% of the total signal. The MACSIS project (Altamura et al., 2023) further revealed that dark matter, galaxy, and gas spins are often misaligned, with the rotational kSZ signal peaking in massive, unrelaxed clusters. Cosmological implications have also been explored: Chluba & Mannheim (2002) examined how rotation impacts the CMB temperature and polarization, while Hamden et al. (2010) used line-of-sight velocity mapping to detect tangential motions and cluster kinematics.
Despite these efforts, most prior studies focus on individual or small sets of clusters, lacking a unified statistical framework. In this work, we address this gap by systematically quantifying spin features across a large sample of galaxy clusters using redshift information from the SDSS DR12 (Alam et al., 2015) spectroscopic catalog. Our analysis explores rotation signatures and their dependence on cluster morphology, dynamical state, and environment, providing one of the first statistical assessments of cluster spin in a large observational dataset.
The structure of this letter is as follows. Section 2 outlines the galaxy group catalog and the selection of member galaxies. Section 3 presents the statistical detection of the cluster spin signal, as well as the connections between cluster spin, central galaxy spin, and the orientation of large-scale filaments. Sections 4 and 5 provide a summary and discussion of our results.
@@ 2 Data and Methodology
This section describes the galaxy and group samples used in this study (Section 2.1) and outlines the method employed to quantify the spin of galaxy clusters (Section 2.2).
@@ 2.1 Galaxy and Group Samples
We use the spectroscopic galaxy catalog presented by Tempel et al. (2017), based on data from the Sloan Digital Sky Survey Data Release 12 (Alam et al., 2015). To ensure data quality, the catalog was cleaned by applying a Petrosian $r$-band magnitude limit of $m_{r}=17.77$ mag, effectively removing spurious or poorly measured entries. The resulting sample consists of 584,449 galaxies, each with group ID assignments, CMB-frame redshifts, and various diagnostic parameters.
@@
The corresponding group catalog was constructed from the same galaxy sample using the Friends-of-Friends (FoF) algorithm (Huchra & Geller, 1982; Zeldovich et al., 1982; Barnes et al., 1985; Dolag et al., 2009). This algorithm links neighboring galaxies based on a defined projected linking length, grouping them into coherent systems. To account for redshift-dependent projection effects, the linking length was scaled with distance following the procedure described in Tempel et al. (2014b). Group memberships were further refined according to the method detailed in Tempel et al. (2016), including a check for candidate merging systems. The final catalog contains 88,662 galaxy groups with at least two member galaxies. Each group is characterized by its richness, Cartesian coordinates, and group ID.
@@
Figure 1: A schematic illustration of the method for identifying the projected rotation axis of a galaxy cluster. Star symbols represent the central galaxy (black) and the member galaxies (red and blue). The dashed line marks the projected rotation axis, which divides the projected plane into two regions: Region A (red) and Region B (blue). If the cluster is rotating in the direction indicated by the blue arrow, member galaxies in Region A are expected to predominantly exhibit red-shifted velocities, while those in Region B are expected to appear mostly blue-shifted.In this illustration, the projected rotation axis is rotated by an angle $\theta_{\rm max}$ from the reference axis (the x-axis, dotted line); see the main text for further details.
@@
Figure 2: Distributions of the basic properties of the galaxy cluster sample used in this study. From left to right, the panels show: (1) cluster mass $\log(M_{\mathrm{cl,200}}/M_{\odot})$; (2) cluster radius $R_{\mathrm{cl,200}}$ (in Mpc); (3) number of member galaxies $N_{\mathrm{gal}}$ (richness); (4) cluster redshift $Z_{\mathrm{cl}}$; and (5) spectroscopic redshift error $Z_{\mathrm{error}}$.
@@ 2.2 Identification of Cluster Spin
in which $Z_{A}$, $Z_{B}$ and $Z_{0}$ are the mean redshifts of galaxies in region A, region B, and all member galaxies of the cluster, respectively. It is worth noting that, according to the above formula, whether $Z_{0}$ is taken as the mean redshift of all member galaxies of the cluster or the redshift of the central galaxy does not affect the value of $\Delta Z_{\rm max}$. The $\Delta Z_{\rm max}$ serves as a proxy for the cluster’s spin amplitude. The value of $\theta_{\rm max}$ is then used to determine the orientation of the rotation axis of the cluster in the projected plane. It is worth noting that, in the process of determining $\theta_{\rm max}$ as shown in Figure 1, we sometimes encounter cases where the same value of $\Delta Z_{\rm max}$ corresponds to multiple values of $\theta_{\rm max}$. This phenomenon is caused by the sparse spatial distribution and the limited number of member galaxies, as discussed in the Appendix A. In addition, the root-mean-square (RMS) deviation of the redshifts of member galaxies, denoted as $Z_{\rm rms}$, is also calculated to quantify the redshift dispersion within each cluster. We further adopt the ratio $Z_{\rm rms}/\Delta Z_{\max}$ as a proxy for the cluster dynamical “temperature”, similar to the filament dynamical temperature defined in Wang et al. (2021).
To assess the statistical significance of the detected spin signal, we perform a Monte Carlo randomization test, which also used in Wang et al. (2021, 2025b); Tang et al. (2025).
@@
  • Re-calculate $\Delta Z_{\rm max}^{\rm rand}$ using the same rotation axis as determined above.
@@
  • Compute the significance of the observed $\Delta Z_{\rm max}$ using:
@@
Following the conventional definition of galaxy clusters, we select systems with total masses $M>10^{14}\,M_{\odot}$ from the group catalog. The resulting sample comprises 2,170 galaxy clusters, which are used for the following analysis of cluster spin and compared against randomized control samples. The basic statistical properties of the full cluster sample, including virial mass $M_{200}$, virial radius $R_{200}$, richness $N_{\mathrm{gal}}$, and redshift $z$, are summarized in Figure 2. In the third panel, we can see that the sample is roughly divided into two equal subsamples at $N_{\mathrm{gal}}=10$. This division will be taken into account in the subsequent analysis. In addition, the last panel shows that the peak of the spectroscopic redshift error is around $2.5\times 10^{-5}$, which means that, based on the method described above, any $\Delta Z_{\rm max}$ less than or equal to $z_{\rm error}$ is not meaningful.
@@ 3 Result
Figure 3: Left: The cumulative distribution of the redshift difference, $\Delta Z_{\rm max}$, between galaxies in region A and region B. The solid red lines in the bottom panel represent the distribution of the measured cluster spin signal, while the 10,000 gray lines correspond to randomized trials. The black solid lines indicate the median values of the random samples. The top panels display the sample-to-sample distance between the real clusters in the observation and the random samples, expressed in units of the standard deviation of the random trials. The upper x-axes show the rotation speed of the filaments, calculated as $\mu=\frac{1}{2}c\times\,{\rm\Delta Z_{Max}}$. Middle: Similar to the left panel, but show clusters with member galaxies more than 10 ($N_{gal}>10$). Right: The significance of the cluster spin signal as a function of the ratio $Z_{\mathrm{rms}}/\Delta Z_{\rm max}$. Each point represents a cluster, colored by the number of member galaxies in regions A and B ($N_{\mathrm{gal}}^{A,B}$). The vertical dashed line marks $Z_{\mathrm{rms}}/\Delta Z_{\rm max}=1$.
@@ 3.1 Global Detection of Cluster Spin Signals
$\mu=\frac{1}{2}c\times\Delta Z_{\rm max},$ (3)
where $c$ is the speed of light and $\Delta Z_{\rm max}$ is the measured maximum redshift difference between the two regions. The observed cumulative distribution exhibits a consistent and significant excess at higher values of $\Delta Z_{\rm max}$ compared to the random samples. As shown in the top sub-panel, this deviation exceeds $150\sigma$ at $\sim 380{\,\rm km\,s^{-1}}$ and remains significant across a wide range of $\Delta Z_{\rm max}$, indicating a statistically robust signature of coherent rotational motion within the observed galaxy clusters.
The middle panel shows the same analysis, but restricted to clusters with more than 10 member galaxies ($N_{\mathrm{gal}}>10$). The results are qualitatively similar, with the observed clusters again displaying a significant excess in $\Delta Z_{\rm max}$ relative to the randomized controls. Notably, the sample-to-sample deviation shown in the upper sub-panel is even more pronounced than in the full sample, with the deviation exceeding $300\sigma$ at $\sim 450~\mathrm{km\,s}^{-1}$, indicating a higher statistical confidence for the rotational signal in richer systems.
The right panel displays the significance of the cluster spin signal as a function of the ratio $Z_{\mathrm{rms}}/\Delta Z_{\rm max}$ for individual clusters, where the x-axis represents the dynamical temperature as introduced in the study of filament spin by Wang et al. (2021). Each point corresponds to a cluster and is colored according to the number of member galaxies in regions A and B ($N_{\mathrm{gal}}^{A,B}$). The vertical dashed line marks $Z_{\mathrm{rms}}/\Delta Z_{\rm max}=1$. Two key trends are evident: (1) clusters with lower dynamical temperatures (i.e., smaller $Z_{\mathrm{rms}}/\Delta Z_{\rm max}$ ratios) tend to exhibit higher significance, indicating that the spin signal is more prominent in dynamically colder systems; and (2) at a fixed dynamical temperature, clusters with a larger number of member galaxies (indicated by the yellow color) show higher significance, suggesting that richer clusters provide a more robust detection of coherent rotation.

### v3

@@ Abstract
The spin of galaxy clusters encodes key information about their formation, dynamics, and the influence of large-scale structure. However, whether clusters possess statistically significant spin and how to measure it observationally remain open questions. Here, we present the first observational, statistical detection of coherent spin in galaxy clusters, using two samples of 2,170 and 1,329 systems with $M>10^{14}\,M_{\odot}$, selected from two publicly available group catalogs (Tempel et al. (2017) and Yang et al. (2012)) constructed with two different algorithms and but both based primarily on SDSS galaxies. Cluster spin is quantified by identifying the orientation in the projected plane that maximizes the redshift difference ($\Delta Z_{\rm max}$) between member galaxies in two regions divided by a trial axis. We find compelling statistical evidence for coherent rotation, as the observed $\Delta Z_{\rm max}$ distribution departs markedly from the randomized controls, exhibiting pronounced deviations near $380\,\mathrm{km\,s^{-1}}$. Stacked visualizations confirm the spatial segregation of redshifted and blueshifted galaxies across the rotation axis. The radial profile of the rotational velocity indicates that it increases as a function of radius. The cluster rotation speed increases with mass, from $\sim 360~\mathrm{km\,s}^{-1}$ at $10^{14}M_{\odot}$ to $\sim 693~\mathrm{km\,s}^{-1}$ at $10^{15}M_{\odot}$. Additionally, cluster spin tends to align parallel with the central galaxy spin and perpendicular to the nearest cosmic filament, particularly in richer systems. These results reveal significant coherent spin in galaxy clusters, shaped by both internal dynamics and large-scale structure.
@@ 1 Introduction
Simulations have also provided valuable insights into cluster spin. Baldi et al. (2017, 2018, 2019) used MUSIC simulations and kinetic Sunyaev-Zel’dovich (kSZ) maps to detect rotational signals averaging 23% of the total signal. The MACSIS project (Altamura et al., 2023) further revealed that dark matter, galaxy, and gas spins are often misaligned, with the rotational kSZ signal peaking in massive, unrelaxed clusters. Cosmological implications have also been explored: Chluba & Mannheim (2002) examined how rotation impacts the CMB temperature and polarization, while Hamden et al. (2010) used line-of-sight velocity mapping to detect tangential motions and cluster kinematics.
Despite these efforts, most prior studies focus on individual or small sets of clusters, lacking a unified statistical framework. In this work, we address this gap by systematically quantifying spin features across a large sample of galaxy clusters using redshift information from the SDSS spectroscopic catalog. Our analysis explores rotation signatures and their dependence on cluster morphology, dynamical state, and environment, providing one of the first statistical assessments of cluster spin in a large observational dataset.
The structure of this letter is as follows. Section 2 outlines the two galaxy group catalogs and the selection of member galaxies. Section 3 presents the statistical detection of the cluster spin signal, as well as the connections between cluster spin, central galaxy spin, and the orientation of large-scale filaments. Sections 4 and 5 provide a summary and discussion of our results.
@@ 2 Data and Methodology
This section is organized as follows. Sections 2.1.1 and 2.1.2 describe the galaxy and group samples used in this study, and Section 2.2 outlines the method employed to quantify the spin of galaxy clusters.
@@ 2.1 Galaxy and Group Samples
To address the potential impact of projection effects and to assess the robustness of our spin measurements, we make use of group catalogs constructed with two different methodologies, as summarized below.
Numerous group catalogs have been constructed from redshift samples obtained by the Sloan Digital Sky Survey (SDSS), using various algorithms such as the C4 algorithm (Miller et al., 2005), friends-of-friends (FoF, Berlind et al., 2006), halo-based group finders (Yang et al., 2007), and density-based methods (Graham & Cappellari, 2023). In particular, FoF-based catalogs are known to be susceptible to projection effects, which may bias measurements of galaxy cluster angular momentum. To minimize the influence of any single group finder and to perform an explicit consistency check, we therefore employ both an FoF-based group catalog as Sample-1 (Tempel et al., 2017) and a halo-based group catalog as Sample-2 (Yang et al., 2012). The latter follows the halo-based method originally developed by Yang et al. (2007), which has been shown to effectively mitigate projection effects. As we demonstrate in the following sections, the spin signals measured from these two independent samples are consistent with each other, supporting the robustness of our results against the choice of group finder.
2.1.1 Sample-1
@@
For our Sample-1, we use the spectroscopic galaxy catalog presented by Tempel et al. (2017), based on data from the Sloan Digital Sky Survey Data Release 12 (Alam et al., 2015). To ensure data quality, the catalog was cleaned by applying a Petrosian $r$-band magnitude limit of $m_{r}=17.77$ mag, effectively removing spurious or poorly measured entries. The resulting sample consists of 584,449 galaxies, each with group ID assignments, CMB-frame redshifts, and various diagnostic parameters.
@@
The corresponding group catalog was constructed from the same galaxy sample using the Friends-of-Friends (FoF) algorithm (Huchra & Geller, 1982; Zeldovich et al., 1982; Barnes et al., 1985; Dolag et al., 2009). This algorithm links neighboring galaxies based on a defined projected linking length, grouping them into coherent systems. To account for redshift-dependent projection effects, the linking length was scaled with distance following the procedure described in Tempel et al. (2014b). Group memberships were further refined according to the method detailed in Tempel et al. (2016), including a check for candidate merging systems, and group centre was calculated as the geometrical centre of all member galaxies without any luminosity or mass weighting as described in Tempel et al. (2016). The final catalog comprises 88,662 galaxy groups containing a total of 287,245 galaxies. Each group is described by its richness, Cartesian coordinates, and group ID.
2.1.2 Sample-2
For our Sample-2, we select clusters from the catalog of Yang et al. (2012), which is constructed from spectroscopic galaxy data in the New York University Value-Added Galaxy Catalog (NYU-VAGC; Blanton et al. 2005), itself based on SDSS DR7 (Abazajian et al., 2009). In the NYU-VAGC-based sample of Yang et al. (2012), galaxies in the Main Galaxy Sample with extinction-corrected apparent magnitude $r<17.72$, redshifts in the range $0.01\leq z\leq 0.20$, and redshift completeness $C_{z}>0.7$ are selected, resulting in a spectroscopic sample of 639,359 galaxies covering 7,748 square degrees.
The corresponding group catalog was constructed from the same galaxy sample using the halo-based group finder of Yang et al. (2007). This group-finding algorithm is iterative and relies on an adaptive filter modeled after the general properties of dark matter halos. In contrast to the traditional FoF method, it is capable of identifying groups with only a single member, and the group center is determined as the luminosity-weighted geometric center of its member galaxies. The final catalog comprises 472,416 galaxy groups containing a total of 639,359 galaxies. Each group is characterized by its celestial coordinates, richness, and group ID.
@@ 2.1.2 Sample-2
Figure 1: A schematic illustration of the method for identifying the projected rotation axis of a galaxy cluster. Star symbols represent the the geometrical centre of all galaxies in the group(black) and the member galaxies (red and blue). The dashed line marks the projected rotation axis, which divides the projected plane into two regions: Region A (red) and Region B (blue). If the cluster is rotating in the direction indicated by the blue arrow, member galaxies in Region A are expected to predominantly exhibit red-shifted velocities, while those in Region B are expected to appear mostly blue-shifted.In this illustration, the projected rotation axis is rotated by an angle $\theta_{\rm max}$ from the reference axis (the x-axis, dotted line); see the main text for further details.
@@
Figure 2: Distributions of the basic properties of the galaxy clusters used in this study for both Sample-1 (blue line) and Sample-2 (orange line). From left to right, the panels show: (1) cluster mass $\log(M_{\mathrm{cl}}/M_{\odot})$; (2) cluster radius $R_{\mathrm{cl}}$ (in Mpc); (3) number of member galaxies $N_{\mathrm{gal}}$ (richness); (4) cluster redshift $Z_{\mathrm{cl}}$; and (5) spectroscopic redshift error $Z_{\mathrm{error}}$.
@@ 2.2 Identification of Cluster Spin
in which $Z_{A}$, $Z_{B}$ and $Z_{0}$ are the mean redshifts of galaxies in region A, region B, and all member galaxies of the cluster, respectively. It is worth noting that, according to the above formula, whether $Z_{0}$ is taken as the mean redshift of all member galaxies of the cluster or the redshift of the central galaxy does not affect the value of $\Delta Z_{\rm max}$. The $\Delta Z_{\rm max}$ serves as a proxy for the cluster’s spin amplitude. The value of $\theta_{\rm max}$ is then used to determine the orientation of the rotation axis of the cluster in the projected plane. It is worth noting that, in the process of determining $\theta_{\rm max}$ as shown in Figure 1, we sometimes encounter cases where the same value of $\Delta Z_{\rm max}$ corresponds to multiple values of $\theta_{\rm max}$. This phenomenon is caused by the sparse spatial distribution and the limited number of member galaxies, as discussed in the Appendix A(corresponding to Sample-1). In addition, the root-mean-square (RMS) deviation of the redshifts of member galaxies, denoted as $Z_{\rm rms}$, is also calculated to quantify the redshift dispersion within each cluster. We further adopt the ratio $Z_{\rm rms}/\Delta Z_{\max}$ as a proxy for the cluster dynamical “temperature”, similar to the filament dynamical temperature defined in Wang et al. (2021).
To assess the statistical significance of the detected spin signal, we perform a Monte Carlo randomization test, which was alsoalso used in Wang et al. (2021, 2025b); Tang et al. (2025).
@@
  • Calculate $\Delta Z_{\rm max}^{\rm rand}$ by, for each Monte Carlo realization, finding the rotation axis that maximizes $\Delta Z_{\rm max}$ and recording that maximum value. the same selection procedure is applied to the randomized samples as to the real cluster, ensuring fair comparison.
@@
  • Compute the significance of each observed cluster $\Delta Z_{\rm max}$ using:
@@
Following the conventional definition of galaxy clusters, we select systems with total masses $M>10^{14}\,M_{\odot}$ from the two group catalogs. The resulting sample comprises 2,170 and 1329 galaxy clusters respectively, which are used for the following analysis of cluster spin and compared against randomized control samples. The basic statistical properties of the cluster samples, including virial mass $M$, virial radius $R$, richness $N_{\mathrm{gal}}$, and redshift $z$, are summarized in Figure 2. The similar distributions of the basic properties of the two samples ensure the feasibility of a comparative analysis, particularly with respect to mass and radius. The slight differences in $N_{\mathrm{gal}}$ and redshift $z$ mainly arise from the different galaxy input catalogs and the distinct group-finding algorithms adopted in the two cases. In the third panel, we can see that for sample-1 is roughly divided into two equal subsamples at $N_{\mathrm{gal}}=10$. This division will be taken into account in the subsequent analysis. The corresponding richness distributions for both samples are also displayed in the left panel of Figure A2 for reference. In addition, the last panel shows that the peak of the spectroscopic redshift error occurs at roughly $2.5\times 10^{-5}$ for both samples, while Sample-2 exhibits a somewhat broader range, with a maximum reaching only $\sim 2\times 10^{-4}$. Part of the difference arises from the known quantization of inverse variance in the SDSS DR7 pipeline, which artificially increases the reported flux uncertainties in low-S/N regions; this issue is largely mitigated in the DR12 reduction, leading to systematically smaller and smoother error estimates. which means that, based on the method described above, any $\Delta Z_{\rm max}$ less than or equal to $z_{\rm error}$ is not meaningful.
@@ 3 Result
Figure 3: The global detection of cluster spin for Sample-1 (top row) and Sample-2 (bottom row). Left: The cumulative distribution of the redshift difference, $\Delta Z_{\rm max}$, between galaxies in region A and region B. The solid red lines in the bottom panel represent the distribution of the measured cluster spin signal, while the 10,000 gray lines correspond to randomized trials. The black solid lines indicate the median values of the random samples. The top panels display the sample-to-sample distance between the real clusters in the observation and the random samples, expressed in units of the standard deviation of the random trials. The upper x-axes show the rotation speed of the filaments, calculated as $\mu=\frac{1}{2}c\times\,{\rm\Delta Z_{Max}}$. Middle: Similar to the left panel, but show clusters with member galaxies more than 10 ($N_{gal}>10$). Right: The significance of the cluster spin signal as a function of the ratio $Z_{\mathrm{rms}}/\Delta Z_{\rm max}$. Each point represents a cluster, colored by the number of member galaxies in regions A and B ($N_{\mathrm{gal}}^{A,B}$). The vertical dashed line marks $Z_{\mathrm{rms}}/\Delta Z_{\rm max}=1$.
@@ 3.1 Global Detection of Cluster Spin Signals
The measured coherent spin signals of galaxy clusters for the two samples are presented in Figure 3, with the top row corresponding to Sample-1 and the bottom row to Sample-2. Qualitatively consistent results are obtained for the two cluster samples defined by different group-finding algorithms. The observed cumulative distributions show a consistent and significant excess at higher values of $\Delta Z_{\rm max}$ compared to the random samples, further indicating that the choice of algorithm does not affect the measurement of the spin signal and validating the robustness and reliability of our method. Some quantitative differences between the two samples are present. For example, the signal reaches its maximum confidence level at $\sim 380\ \mathrm{km\,s^{-1}}$ for Sample-1 and at $\sim 300\ \mathrm{km\,s^{-1}}$ for Sample-2, which we attribute to differences in the underlying galaxy samples and in the cluster-finding algorithms. In the following, we focus on the results from Sample-1 for illustrative purposes, noting that Sample-2 yields consistent trends with slightly different amplitudes.
@@
$\mu=\frac{1}{2}c\times\frac{\Delta Z_{\rm max}}{1+Z_{\rm cl}},$ (3)
where $c$ is the speed of light and $\Delta Z_{\rm max}$ is the measured maximum redshift difference between the two regions, $Z_{\mathrm{cl}}$ is the cluster redshift, calculated as an average over all cluster members. The observed cumulative distribution exhibits a consistent and significant excess at higher values of $\Delta Z_{\rm max}$ compared to the random samples. As shown in the top sub-panel, the cumulative deviation reaches a level corresponding to an apparent $\sim 100\sigma$ when integrated over all clusters (around $\sim 380\,\mathrm{km\,s^{-1}}$). This value reflects the aggregated, sample-to-sample confidence rather than the significance of any individual cluster. The deviation remains substantial across a wide range of $\Delta Z_{\rm max}$, indicating a statistically robust collective trend of coherent rotational motion among the observed clusters. For completeness, the probability density distribution (PDF) of $\Delta Z_{\rm max}$ is displayed in the middle panel of Figure A2, illustrating the detailed shape of the distribution.
The middle panel shows the same analysis, but restricted to clusters with more than 10 member galaxies ($N_{\mathrm{gal}}>10$). The results are qualitatively similar, with the observed clusters again displaying a significant excess in $\Delta Z_{\rm max}$ relative to the randomized controls. Notably, the cumulative sample-to-sample deviation shown in the upper sub-panel is even more pronounced than in the full sample, reaching an apparent level of $\sim 120\sigma$ at $\sim 379\,\mathrm{km\,s^{-1}}$. This value represents an aggregated confidence across richer clusters rather than the significance of individual systems, indicating a stronger collective rotational tendency in these high-richness clusters. The right panel displays the significance of the cluster spin signal as a function of the ratio $Z_{\mathrm{rms}}/\Delta Z_{\rm max}$ for individual clusters, where the x-axis represents the dynamical temperature as introduced in the study of filament spin by Wang et al. (2021). Each point corresponds to a cluster and is colored according to the number of member galaxies in regions A and B ($N_{\mathrm{gal}}^{A,B}$). The vertical dashed line marks $Z_{\mathrm{rms}}/\Delta Z_{\rm max}=1$. Two key trends are evident: (1) clusters with lower dynamical temperatures (i.e., smaller $Z_{\mathrm{rms}}/\Delta Z_{\rm max}$ ratios) tend to exhibit higher significance, indicating that the spin signal is more prominent in dynamically colder systems; and (2) at a fixed dynamical temperature, clusters with a larger number of member galaxies (indicated by the yellow color) show higher significance, suggesting that richer clusters provide a more robust detection of coherent rotation.
We also performed Kolmogorov–Smirnov (KS) tests to quantify the differences between the observed sample and the randomized realizations. For the direct two-sample comparison, we obtained a KS statistic of $D=0.6986$ with a $p$-value $p\approx 0.0\mathrm{e}{+}00$ (i.e., $p\ll 10^{-300}$). For the comparison of the corresponding cumulative distributions (the red and black curves in Figure 3), we found $D=0.305$ with $p=1.87\times 10^{-41}$. These extremely small $p$-values indicate that the observed and randomized samples are statistically inconsistent, consistent with the clear visual separation between the red and black curves in Figure 3.